OpenStudy (liizzyliizz):

for logx+log(x-3)=1 is ( x =-2) and (x=5 )?

myininaya (myininaya):

\[\log(x(x-3))=1\] \[10^{1}=x(x-3) =>10=x^2-3x=>0=x^2-3x-10\] \[(x-5)(x+2)=0=>x=5 or x=-2\] but x=-2 will not work

OpenStudy (liizzyliizz):

thank you :)